Best Rentals charges a daily fee plus a mileage fee for renting its cars. Barney was charged $135.00 for 3 days and 300 miles, while Mary was charged $250.00 for 5 days and 600 miles. What does Best Rentals charge per day and per mile?
Answer by mananth(16946)

Barney was charged $135.00 for 3 days and 300 miles, while Mary was charged $250.00 for 5 days and 600 miles. What does Best Rentals charge per day and per mile?
let charge per day be $ x
& charge per mile be $ y
3x+300y =135 ( nultiply by 5 )
5x+600y= 250 ( multiply by 3) To eliminate x
3(5x+600y)-5(3x+300y) = 3*250 - 5*135
15x+1800y-15x-1500y= 750-675
300y = 75
y= 75/300
y= 1/4 = 25 cents per mile

3x+300 *1/4 =135
3x+75=135
3x=135-75
3x=60
x= 20
The charges are $20 per day and 25 cents per mile